OpenStudy (anonymous):

2. Suppose that y varies inversely with x, and y = 0.2 when x = 2. What is an equation for the inverse variation? (1 point)

jimthompson5910 (jim_thompson5910):

y = k/x 0.2 = k/2 0.2*2 = k 0.4 = k k = 0.4 So the equation is \[\Large y = \frac{0.4}{x}\]